### v2

# Bibliography.bib

Abstract.
In this paper, we study how the cohomology of nilpotent groups is affected by Lipschitz maps. We show that, given a smooth Lipschitz map $f$ between two simply-connected nilpotent Lie groups $G$ and $H$, there is a map $\psi$ that induces an ergodic measure on the space of functions from $G$ to $H$. We call such maps ergodic maps.
We show that when $\psi$ is an ergodic map, the pullback $\psi^{*}\omega$ of a differential form $\omega$ admits a well-defined amenable average $\overline{\psi^{*}}\omega$, and $\overline{\psi^{*}}$ is a homomorphism of cohomology algebras. In the case that $f$ is a quasi-isometry, the ergodic map $\psi$ is also a quasi-isometry, and $\overline{\psi^{*}}$ is an isomorphism. This lets us generalize and provide a simplified, self-contained proof of the theorem due to Shalom, Sauer, and Gotfredsen–Kyed that quasi-isometric nilpotent groups have isomorphic cohomology algebras.
1. Introduction
The question of when two nilpotent groups are quasi-isometric has motivated many developments in geometric group theory. Since any finitely generated nilpotent group is a cocompact lattice in a nilpotent Lie group, it suffices to consider nilpotent Lie groups. Clearly, two isomorphic nilpotent Lie groups are quasi-isometric, but it remains open whether two quasi-isometric groups are isomorphic.
One of the first results in this direction was Pansu’s theorem [PansuCroissance, Pansu], which says that the asymptotic cone of a nilpotent group is isomorphic to its associated graded group, and that two quasi-isometric nilpotent groups must have isomorphic asymptotic cones. Shalom [Shalom] was the first to exhibit two nilpotent groups with the same associated graded group which are not quasi-isometric, by showing that the Betti numbers of finitely-generated nilpotent groups are invariant under quasi-isometry. Sauer [Sauer] extended this result, showing that the cup product structure is likewise preserved under quasi-isometry, and Gotfredsen and Kyed extended this to arbitrary nilpotent Lie groups [GotfredsenKyed]. More recently, Llosa Isenrich, Pallier, and Tessera [LlosaIsenrichPallierTessera] gave an example of two groups with isomorphic asymptotic cones which can be distinguished using their Dehn functions.
Shalom and Sauer’s results are particularly promising because they point to a possible strategy for solving the conjecture. The cup product structure is the first of a series of higher cohomology operations that operate on two, three, or more cohomology classes at a time. If these higher operations are likewise invariant under quasi-isometry, in the sense that a quasi-isometry from $G$ to $H$ induces a $C_{\infty}$–morphism from $H^{*}(H)$ to $H^{*}(G)$, that would imply that the minimal models of $H^{*}(H)$ and $H^{*}(G)$ are isomorphic [KadeCohom, Theorem 9.1], and in turn, that $G$ and $H$ are isomorphic.
In this paper, we generalize the result of Shalom, Sauer, and Gotfredsen–Kyed to maps that are not quasi-isometries and give a new, more geometric proof of their results based on pullbacks of differential forms. Before stating our main theorems, we introduce some notation. Let $\phi\colon G\to H$ be a Lipschitz map between two amenable connected Riemannian Lie groups, and let $\omega$ be a left-invariant differential form with complex coefficients. Then the pullback $\phi^{*}\omega$ is a (possibly non-smooth) bounded differential form on $G$, and since $G$ is amenable, one can average $\phi^{*}\omega$ to obtain a left-invariant differential form.
That is, let $L^{\infty}(G)$ be the set of complex-valued $L^{\infty}$ functions on $G$, let $\mathfrak{g}$ be the Lie algebra of $G$, and let $M\colon L^{\infty}(G)\to\mathbb{C}$ be a left-invariant mean. For every left-invariant $d$–vector field $\lambda\in\wedge^{d}\mathfrak{g}$ on $G$, we can evaluate $\phi^{*}\omega$ on $\lambda$ to obtain a bounded function $\phi^{*}\omega(\lambda)$ on $G$, and there is a unique left-invariant form $\overline{\phi^{*}}\omega\in C^{d}(\mathfrak{g})$ such that
$\overline{\phi^{*}}\omega(\lambda):=M(\phi^{*}\omega(\lambda)),\qquad\text{ for all }\lambda\in\wedge^{d}\mathfrak{g}.$ (1)
We denote this by $\overline{\phi^{*}}\omega$ because it is the result of averaging the pullback $\phi^{*}\omega$.
It is hard to characterize $\overline{\phi^{*}}$ in the general case, because a Lipschitz map from $G$ to $H$ can have different behavior at different locations and scales. In this paper, however, we will show two main results. First, if $\phi$ is an ergodic map (i.e., it satisfies an ergodicity condition that we will define below), then $\overline{\phi^{*}}$ is a homomorphism of cohomology algebras. Second, for any map $f$ from $G$ to $H$, there is a limit of translates of $f$ which is an ergodic map. In particular, if $f$ is a quasi-isometry from $G$ to $H$, then there is an ergodic quasi-isometry $\phi$, and $\overline{\phi^{*}}$ is an isomorphism. This lets us recover the theorems of Shalom, Sauer, and Gotfredsen–Kyed and extend their results from quasi-isometries to the broader class of Lipschitz maps between nilpotent groups.
Before we state our results, we first define ergodic maps (see Section 3 for full details). Let $Y_{0}$ be the space of smooth maps $f$ from $G$ to $H$ such that all derivatives of $f$ are bounded, and $f(\operatorname{id}_{G})=\operatorname{id}_{H}$. We can define a right action of $G$ on $Y_{0}$ by letting $(f\cdot g)(x):=f(g)^{-1}f(g\cdot x)$. Since $(f\cdot g)(\operatorname{id}_{G})=\operatorname{id}_{H}$ and $f\cdot g$ has bounded derivatives, the closure of the orbit $f\cdot G$ is compact by the Arzelà–Ascoli theorem.
For any $n>0$, let $B_{n}\subset G$ be the ball of radius $n$ centered at $\operatorname{id}_{G}$. Given $f\in Y_{0}$, there is a measure $\mu_{n}$ on $Y_{0}$ such that for any continuous function with compact support $A\colon Y_{0}\to\mathbb{R}$,
$\int A\,\mathrm{d}\mu_{n}:=\fint_{B_{n}}A(f\cdot g)\,\mathrm{d}g,$
i.e., $\int A\,\mathrm{d}\mu_{n}$ is the expected value of $A(f\cdot g)$ for $g$ drawn uniformly from $B_{n}$, see Section 2 for notation. We say that $f$ is an ergodic map if the measures $\mu_{n}$ converge weakly to some limit measure $\mu$, and $\mu$ is an ergodic measure on $Y_{0}$ with respect to the right action of $G$ defined above.
For example, the map $f_{1}\colon\mathbb{R}\to\mathbb{R}^{2}$, $f_{1}(x)=(x,\sin x)$ is ergodic, with limit measure $\mu$ supported on maps of the form
$h_{\theta}(x)=(x,\sin(x+\theta)-\sin\theta),$
but $f_{2}(x)=(x,|x|)$ is not; its limit measure is a two-point measure supported on $x\mapsto(x,\pm x)$.
Our main result states that any ergodic map induces an algebra homomorphism from $H^{*}(\mathfrak{h})$ to $H^{*}(\mathfrak{g})$.
Theorem 1.1.
Let $G$ and $H$ be simply connected nilpotent Lie groups, and let $\psi\in Y_{0}$ be an ergodic map. For any $d\geq 0$, any left-invariant form $\omega\in C^{d}(\mathfrak{h})$, and any left-invariant $d$–vector field $\lambda\in\wedge^{d}\mathfrak{g}$, let
$\overline{\psi^{*}}\omega(\lambda):=\lim_{n\to\infty}\fint_{B_{n}}\psi^{*}\omega(\lambda)\,\mathrm{d}g.$ (2)
Then this limit exists, and $\overline{\psi^{*}}$ induces an homomorphism $\overline{\psi^{*}}\colon H^{*}(\mathfrak{h})\to H^{*}(\mathfrak{g})$ of cohomology algebras over $\mathbb{C}$. If $\psi$ is a quasi-isometry, then $\overline{\psi^{*}}$ is an isomorphism.
Furthermore, ergodic maps are common, as shown by the following result.
Proposition 1.2.
Let $G$ and $H$ be simply connected nilpotent Lie groups, and let $f\in Y_{0}$. Then there is a sequence $g_{n}\in G$ such that $f\cdot g_{n}$ converges to a map $\phi\in Y_{0}$ uniformly on compact subsets, and $\phi$ is ergodic. In particular, if $f$ is a quasi-isometry or a quasi-isometric embedding, then so is $\phi$.
As a consequence of Theorem 1.1 and Proposition 1.2, we prove the following theorem.
Theorem 1.3.
Let $G$ and $H$ be quasi-isometric simply connected nilpotent Lie groups. Then there are a smooth Lipschitz quasi-isometry $\phi\colon G\to H$ and a left-invariant mean $M\colon L^{\infty}(G)\to\mathbb{C}$ such that $\overline{\phi^{*}}\colon H^{*}(\mathfrak{h})\to H^{*}(\mathfrak{g})$ is an isomorphism of cohomology algebras (over $\mathbb{C}$).
This implies the results of Shalom, Sauer, and Gotfredsen–Kyed and constructs an explicit isomorphism from $H^{*}(\mathfrak{h})$ to $H^{*}(\mathfrak{g})$. If one could show that this isomorphism also preserves higher cohomology operations, it would imply that $\mathfrak{g}$ and $\mathfrak{h}$ are isomorphic, but this remains open.
Since Theorem 1.1 is not limited to quasi-isometries, it can be applied to other maps. One example, which we learned from Kyle Hansen, is the case of maps with positive asymptotic degree (see [BerdGuthManin]). If $G$ and $H$ are simply connected nilpotent groups of the same dimension, $\psi\colon G\to H$ is Lipschitz, and $\omega$ is the volume form on $H$, we say that $\psi$ has positive asymptotic degree if
$\limsup_{R\to\infty}\frac{\int_{B^{G}_{R}}\psi^{*}\omega}{\big{|}B^{G}_{R}\big{|}}>0,$
where $B^{G}_{R}$ is the ball of radius $R$ around the origin in $G$. If $\psi$ is an ergodic map with positive asymptotic degree, then $\overline{\psi^{*}}\omega\neq 0$, so by Poincaré duality (see the proof of Theorem 1.1), $\overline{\psi^{*}}$ is injective on cohomology. Hence, for example, there is no ergodic Lipschitz map $\psi$ from the Heisenberg group to $\mathbb{R}^{3}$ with positive asymptotic degree.
Outline of paper
In order to use results on cohomology with coefficients in a unitary representation, we use complex coefficients throughout this paper, so $H^{*}(\mathfrak{g})$ denotes $H^{*}(\mathfrak{g};\mathbb{C})$; by the Universal Coefficient Theorem, our results also hold over real coefficients.
In Section 2 we start by defining the locally compact space $Y_{0}$ of smooth maps with bounded derivatives from $G$ to $H$, and an action of $G$ on $Y_{0}$. We prove Proposition 1.2 in Section 3 by showing that if $\phi\in Y_{0}$, then its limit measure is a $G$–invariant probability measure. We then use the Krein–Milman theorem to construct an ergodic measure $\mu$, and use an ergodic theorem of E. Lindenstrauss (see Theorem 3.2) to construct an ergodic $\phi$.
For any left-invariant differential form $\omega\in C^{d}(\mathfrak{h})$ and any $\psi\in Y_{0}$, the pullback $\psi^{*}\omega$ is a smooth bounded $d$–form which depends on $\psi$. In Section 4 we define an algebra $C^{\infty,\infty}(Y_{0})$ of complex-valued functions on $Y_{0}$ such that for each $\alpha\in C^{\infty,\infty}(Y_{0})$ and $\psi\in Y_{0}$, there is a smooth function $\alpha_{\psi}\in C^{\infty}(G)$ with bounded derivatives.
Using this algebra, we construct a cochain complex $C^{*}(\mathfrak{g};C^{\infty,\infty}(Y_{0}))$ that represents forms on $G$ that depend on an element of $Y_{0}$, i.e., for every $\beta\in C^{d}(\mathfrak{g};C^{\infty,\infty}(Y_{0}))$ and $\psi\in Y_{0}$, there is a smooth bounded $d$–form $\beta_{\psi}$ on $G$. This lets us construct a universal pullback map $T^{\sharp}\colon C^{*}(\mathfrak{h})\to C^{*}(\mathfrak{g};C^{\infty,\infty}(Y_{0}))$ such that for each left-invariant form $\omega\in C^{d}(\mathfrak{h})$, and $\psi\in Y_{0}$,
$(T^{\sharp}\omega)_{\psi}=\psi^{*}\omega.$
Such $T^{\sharp}$ is a map of differential graded algebras, and thus induces an algebra homomorphism from the cohomology $H^{*}(\mathfrak{h})$ to $H^{*}(\mathfrak{g};C^{\infty,\infty}(Y_{0}))$, see Lemma 4.1.
In Section 5 we prove a consequence of a vanishing result of Blanc, see Lemma 5.2. We use this result to show that if $\mu$ is ergodic under the action of $G$, $L^{2}(\mu)$ is the set of $L^{2}$ functions on $Y_{0}$, and $M\colon L^{2}(\mu)\to\mathbb{C}$ is the map $M(f):=\int_{Y_{0}}f\,\mathrm{d}\mu$, then $M$ induces an algebra homomorphism from the cohomology ${H}^{*}(\mathfrak{g};C^{\infty,\infty}(Y_{0}))$ to $H^{*}(\mathfrak{g})$, see Lemma 6.2 in Section 6. Since we can factor $\overline{\phi^{*}}=M\circ T^{\sharp}$ (see Lemma 6.1), this implies that $\overline{\phi^{*}}$ is an algebra homomorphism, and thus settles the first part of Theorem 1.1.
Finally, in Section 7, we prove the last parts of Theorem 1.3, and Theorem 1.1. We prove that if $\phi$ is an ergodic quasi-isometry and $\omega$ is the volume form on $H$, then $\overline{\phi^{*}}(\omega)\neq 0$, see Corollary 7.2. Thus, by Poincaré duality, $\overline{\phi^{*}}$ is an injective homomorphism from $H^{*}(\mathfrak{h})$ to $H^{*}(\mathfrak{g})$. By applying the same argument with $G$ and $H$ switched, we obtain an injective homomorphism from $H^{*}(\mathfrak{g})$ to $H^{*}(\mathfrak{h})$, so $H^{*}(\mathfrak{h})$ and $H^{*}(\mathfrak{g})$ are isomorphic algebras, as desired.
Acknowledgments
We thank Roman Sauer for having shared with us his notes with Arthur Bartels on $C_{\infty}$ algebras and morphisms. We also thank Kyle Hansen, Bruce Kleiner, Enrico Le Donne, Fedya Manin, Assaf Naor, and Stefan Wenger for useful discussions and encouragements during various stages of this project. G.A. acknowledges the financial support of the Courant Institute and the AMS-Simons Travel Grant. R.Y. was supported by the National Science Foundation under Grant No. 2005609.
2. Preliminaries
Let $G$ and $H$ be simply connected nilpotent Lie groups. We denote the identities of $G$ and $H$ by $\operatorname{id}_{G}$ and $\operatorname{id}_{H}$. Let $\mathfrak{g}$ and $\mathfrak{h}$ be their Lie algebras; we identify $\mathfrak{g}$ with the tangent space $T_{\operatorname{id}_{G}}G$ and with the space $\mathcal{L}_{G}$ of left-invariant vector fields on $G$.
We fix Haar measures on $G$ and $H$. We denote the Haar measure of a subset $U\subset G$ by $|U|$ and the integral with respect to Haar measure by $\int_{G}\beta(g)\,\mathrm{d}g$. We denote the average with respect to Haar measure by
$\fint_{U}\beta\,\mathrm{d}g:=\frac{1}{|U|}\int_{U}\beta\,\mathrm{d}g.$
We will need to define some spaces of smooth maps with bounded derivatives. Let $C^{\infty,\infty}(G)\subset C^{\infty}(G)$ be the space of smooth functions $\alpha\colon G\to\mathbb{C}$ which have bounded derivatives in the sense that $\mathcal{D}\alpha\in L^{\infty}(G)$ for any left-invariant differential operator $\mathcal{D}$. That is, if $V_{1},\dots,V_{m}\in\mathcal{L}_{G}$ is a basis of the Lie algebra of $G$, then
$V_{i_{1}}\dots V_{i_{d}}\alpha\in L^{\infty}(G)$
for any multi-index $I=(i_{1},\dots,i_{d})$.
Let $U_{1},\dots,U_{m}\in\mathfrak{h}$ be a basis of the Lie algebra of $H$. Suppose that $\alpha\in C^{\infty}(G,H)$. Let $m_{ij}\colon G\to\mathbb{R}$ be the matrix coefficients of $D_{g}\alpha\colon T_{g}G\to T_{\alpha(g)}H$ , i.e,
$D_{g}\alpha(V_{i}(g))=\sum_{j}m_{ij}(g)U_{j}(\alpha(g)).$
We say that $\alpha\in C^{\infty,\infty}(G,H)$ if $m_{ij}\in C^{\infty,\infty}(G)$ for all $i$ and $j$; if so, we say that $\alpha$ has bounded derivatives. Let
$Y_{0}(G,H):=\{\alpha\in C^{\infty,\infty}(G,H):\alpha(\operatorname{id}_{G})=\operatorname{id}_{H}\}.$ (3)
Note that $Y_{0}(G,H)\subset\operatorname{Lip}(G,H)$. We equip $C^{\infty,\infty}(G)$ and $Y_{0}(G,H)$ with the Whitney topology, so that a sequence $\alpha_{1},\alpha_{2},\dots$ converges to $\alpha$ if and only if for every compact $K\subset G$, and every $n\in\mathbb{N}$, we have $\alpha_{i}\to\alpha$ in the $C^{n}$–topology on $K$ for all $n$. This makes $C^{\infty,\infty}(G)$ and $Y_{0}(G,H)$ locally compact Hausdorff spaces. In fact, $C^{\infty,\infty}(G)$ and $Y_{0}(G,H)$ are metrizable. Let $\mathcal{D}_{1},\mathcal{D}_{2},\dots$ be a countable basis of the set of left-invariant differential operators, and $K_{1}\subset K_{2}\subset\dots$ be an exhaustion of compact sets in $G$. For $\alpha,\beta\in C^{\infty,\infty}(G)$, and $i,j\geq 1$, define $p_{i,j}(\alpha,\beta):=\|\mathcal{D}_{i}[\beta-\alpha]\|_{L^{\infty}(K_{j})}$. Then, re-indexing $\{p_{i,j}\}_{i,j\geq 1}=\{p_{k}\}_{k\geq 1}$, we have that
$d_{\infty,\infty}(\alpha,\beta)=\sum_{k=1}^{\infty}\min\left\{p_{k}(\alpha,\beta),2^{-k}\right\},$
is a metric on $C^{\infty,\infty}(G)$. A similar formula defines a metric on $Y_{0}(G,H)$.
3. Ergodic maps
3.1. Definitions
Let $G$ and $H$ be simply connected nilpotent Lie groups and let $Y_{0}=Y_{0}(G,H)$ be as in Section 2. For $g\in G$, and $h\in H$, let $l_{g}\colon G\to G$ and $l_{h}\colon H\to H$ be the left-translation maps.
For any $g\in G$ and $\alpha\in Y_{0}$, let
$\alpha\cdot g:=l_{\alpha(g)}^{-1}\circ\alpha\circ l_{g}.$ (4)
One can check that $\alpha\cdot g\in Y_{0}$ and that for $g_{1},g_{2}\in G$,
$(\alpha\cdot g_{1})\cdot g_{2}=l_{(\alpha\cdot g_{1})(g_{2})}^{-1}\circ l_{\alpha(g_{1})}^{-1}\circ\alpha\circ l_{g_{1}g_{2}}\\ =l_{\alpha(g_{1})^{-1}\alpha(g_{1}g_{2})}^{-1}\circ l_{\alpha(g_{1})}^{-1}\circ\alpha\circ l_{g_{1}g_{2}}=\alpha\cdot(g_{1}g_{2}).$
That is, this is a right action of $G$ on $Y_{0}$. Note that the derivatives of $\alpha\cdot g$ satisfy bounds that depend only on $\alpha$, so for any $\alpha\in Y_{0}$, the closure $\operatorname{clos}(\alpha\cdot G)$ is compact by the Arzelà–Ascoli Theorem.
Definition 3.1.
Let $y\in Y_{0}$. For every $n>0$, there is a measure $\mu_{n}$ on $Y_{0}$ such that for any continuous function $A\colon Y_{0}\to\mathbb{R}$,
$\int A\,\mathrm{d}\mu_{n}:=\fint_{B_{n}}A(y\cdot g)\,\mathrm{d}g.$ (5)
If the $\mu_{n}$’s converge weakly to a measure $\mu$ on $Y_{0}$, we call $\mu$ the limit measure of $y$. Since the orbit $y\cdot G$ has compact closure, if $\mu$ exists, then $\mu$ is a probability measure supported on a compact set, and the amenability of $G$ implies that $\mu$ is $G$–invariant. We say that $y$ is an ergodic map if $\mu$ exists and is ergodic with respect to the action of $G$.
3.2. Constructing ergodic maps
We construct ergodic maps using the Krein–Milman theorem and the following theorem of E. Lindenstrauss [Lindenstrauss, Theorem 1.2].
Theorem 3.2 (Lindenstrauss).
Let $G$ be an amenable group equipped with a left-invariant Haar measure and let $(X,\eta)$ be a probability space equipped with an ergodic left action of $G$. Let $F_{n}\subset G$ be a tempered Følner sequence. Then for any $f\in L_{1}(\eta)$, the following holds for $\eta$–a.e. $x\in X$,
$\lim_{n\to\infty}\fint_{F_{n}}f(g\cdot x)\,\mathrm{d}g=\int_{X}f\,\mathrm{d}\eta.$
A tempered Følner sequence $F_{n}$ is a Følner sequence for $G$ such that there is some $C>0$ such that for all $n$,
$\left|\bigcup_{k\leq n}F_{k}^{-1}F_{n+1}\right|\leq C|F_{n+1}|.$ (6)
In all our uses of Theorem 3.2, we will take $F_{n}=B_{n}(\operatorname{id}_{G})$, which satisfies (6). Throughout the paper we will abbreviate $B_{n}:=B_{n}(\operatorname{id}_{G})$.
To prove Proposition 1.2, we first construct an invariant ergodic probability measure on $Y_{0}$.
Lemma 3.3.
Let $f\in Y_{0}$. Then there is a $G$–invariant ergodic probability measure $\mu$ on $Y_{0}$ with support contained in the closure $K=\operatorname{clos}(f\cdot G)$.
Proof.
We first construct an invariant probability measure $\eta$ supported on $K$. Let $C_{c}(Y_{0})$ be the set of continuous, compactly-supported, complex-valued functions on $Y_{0}$. Let $\operatorname*{\overline{\lim}}$ be an ultralimit, and for any $\omega\in C_{c}(Y_{0})$, let
$\alpha(\omega):=\operatorname*{\overline{\lim}}_{n\to\infty}\fint_{B_{n}}\omega(f\cdot g)\,\mathrm{d}g.$
By the amenability of $G$, $\alpha$ is $G$–invariant. By the Riesz representation theorem, since $Y_{0}$ is locally compact and Hausdorff, there is a $G$–invariant Radon measure $\eta$ on $Y_{0}$ such that $\alpha(\omega)=\int_{Y_{0}}\omega\,\mathrm{d}\eta$ for all $\omega\in C_{c}(Y_{0})$.
Since $K$ is compact and $Y_{0}$ is metrizable, there is a sequence of functions $k_{i}\in C_{c}(Y_{0})$ such that $k_{i}=1$ on $K$ and $\operatorname{supp}(k_{i})\to K$; then $\alpha(k_{i})=1$ for all $i$, so $\eta(K)=1$. Furthermore, if $\omega\in C_{c}(Y_{0})$ has support disjoint from $K$, then $\alpha(\omega)=0$, so $\eta$ is a probability measure supported on $K$.
We can view the set $P$ of $G$–invariant probability measures supported on $K$ as a convex subset of the dual space $C(K)^{*}$, equipped with the weak-* topology. Since $K$ is compact, this set is compact; since it contains $\eta$, it is nonempty. By the Krein–Milman Theorem, the set $E(P)$ of extreme points of $P$ is nonempty; let $\mu\in E(P)$. By extremality, $\mu$ is an ergodic probability measure supported on $K$, as desired. ∎
We use this lemma to prove Proposition 1.2.
Proof of Proposition 1.2.
Let $f\in Y_{0}$, let $K=\operatorname{clos}(f\cdot G)$, and let $\mu$ be a $G$–invariant ergodic probability measure on $K$ as in Lemma 3.3.
Since $Y_{0}$ is metrizable, there is a countable dense set $S\subset C(K)$. For every $s\in S$, Theorem 3.2 implies that there is a $K_{s}\subset K$ such that $\mu(K\setminus K_{s})=0$ and such that for any $y\in K_{s}$,
$\lim_{n\to\infty}\fint_{B_{n}}s(y\cdot g^{-1})\,\mathrm{d}g=\lim_{n\to\infty}\fint_{B_{n}}s(y\cdot g)\,\mathrm{d}g=\int_{Y_{0}}s\,\mathrm{d}\mu.$ (7)
Since $S$ is countable, $\bigcap_{s}K_{s}$ is nonempty; let $\phi\in\bigcap_{s}K_{s}$. Then for any $A\in C(K)$,
$\lim_{n\to\infty}\fint_{B_{n}}A(\phi\cdot g)\,\mathrm{d}g=\int_{Y_{0}}A\,\mathrm{d}\mu,$ (8)
i.e., $\phi$ is ergodic with limit measure $\mu$.
Furthermore, since $\phi\in K=\operatorname{clos}(f\cdot G)$, there is a sequence $g_{1},g_{2},\dots$ such that $f\cdot g_{n}$ converges to $\phi$ on compact subsets, as desired. ∎
4. The universal pullback map
For any left-invariant form $\omega\in\bigwedge^{*}\mathfrak{h}$, the pullback $\phi^{*}\omega$ is a differential form on $G$ which is bounded, but generally not left-invariant.
One of the main ideas of the proof of Theorem 1.3 is that we can write such pullbacks in terms of a space of functions $A:=C^{\infty,\infty}(Y_{0})\subset C(Y_{0})$ (where $C(Y_{0})$ is the set of continuous functions from $Y_{0}$ to $\mathbb{C}$), and the corresponding differential graded algebra $C^{*}(\mathfrak{g};A)$.
Each element $\alpha\in A$ will correspond to a $G$–equivariant operator that sends an element $\phi\in Y_{0}$ to a smooth function $\alpha_{\phi}\in C^{\infty,\infty}(G)$, and each $\beta\in C^{*}(\mathfrak{g};A)$ will correspond to a $G$–equivariant operator that sends an element $\phi\in Y_{0}$ to a differential form $\beta_{\phi}$. We will use this correspondence to construct the following universal pullback map.
Lemma 4.1.
There is a morphism of differential graded algebras $T^{\sharp}\colon C^{*}(\mathfrak{h})\to C^{*}(\mathfrak{g};A)$ such that for every $\omega\in C^{k}(\mathfrak{h})$ and $\phi\in Y_{0}$,
$(T^{\sharp}\omega)_{\phi}=\phi^{*}\omega.$ (9)
Then, in the next Section 5, we will give conditions for $T^{\sharp}\omega$ to be a limit of coboundaries, i.e., a reduced coboundary.
4.1. Equivariant operators and $C^{\infty,\infty}(Y_{0})$
Let $C(Y_{0})$ and $C(G)$ be the spaces of complex-valued continuous functions on $Y_{0}$ and $G$ respectively, equipped with the topology of uniform convergence on compact sets. If $\mathcal{D}\colon Y_{0}\to C(G)$, we say that $\mathcal{D}$ is $G$–equivariant if for all $g,h\in G$ and $\phi\in Y_{0}$,
$\mathcal{D}[\phi\cdot g](h)=\mathcal{D}[\phi](gh),$ (10)
i.e., $\mathcal{D}[\phi\cdot g]=\mathcal{D}[\phi]\circ l_{g}$. Then $\mathcal{D}[\phi]$ has the same symmetries as $\phi$; if $g\in G$ and $\phi\cdot g=\phi$, then $\mathcal{D}[\phi]\circ l_{g}=\mathcal{D}[\phi]$.
We define $\mathsf{Op}^{\infty,\infty}\subset\mathsf{Op}$ to be the set of smooth $G$–equivariant operators, i.e., operators $\mathcal{D}\colon Y_{0}\to C^{\infty,\infty}(G)$ such that $\mathcal{D}$ is continuous with respect to the topology on $C^{\infty,\infty}(G)$ and that satisfy (10).
The most important elements of $\mathsf{Op}^{\infty,\infty}$ in this paper correspond to pullbacks of left-invariant forms on $H$. For example, if $V\in\wedge^{d}\mathfrak{g}$ is a left-invariant field of $d$–vectors on $G$ and $\omega\in C^{d}(\mathfrak{h})$ is a left-invariant $d$–form, then
$\mathcal{D}[\phi]:=\phi^{*}\omega(V)=\omega(\phi_{*}(V))$ (11)
is smooth and satisfies (10), so $\mathcal{D}\in\mathsf{Op}^{\infty,\infty}$.
We will use $\mathsf{Op}^{\infty,\infty}$ to define $C^{\infty,\infty}(Y_{0})$, an algebra of functions on $Y_{0}$. First, we can represent elements of $\mathsf{Op}$ as continuous functions on $Y_{0}$.
Lemma 4.2.
Let $\mathsf{Op}$ be the set of continuous $G$–equivariant operators $\mathcal{D}\colon Y_{0}\to C(G)$. Let $P\colon\mathsf{Op}\to C(Y_{0})$,
$P(\mathcal{D})(\phi):=\mathcal{D}[\phi](\operatorname{id}_{G}).$
Then $P$ is a bijection with inverse $P^{-1}(\alpha)=\mathcal{D}_{\alpha}$, where for any $\alpha\in C(Y_{0})$, we let $\mathcal{D}_{\alpha}\colon Y_{0}\to C(G)$ be the map
$\mathcal{D}_{\alpha}[\phi](h):=\alpha(\phi\cdot h),\text{\qquad}\phi\in Y_{0},h\in G.$ (12)
Proof.
For any $g,h\in G$, $\alpha\in C(Y_{0})$, and $\phi\in Y_{0}$,
$\mathcal{D}_{\alpha}[\phi\cdot g](h)=\alpha(\phi\cdot gh)=\mathcal{D}_{\alpha}[\phi](gh),$
so $\mathcal{D}_{\alpha}\in\mathsf{Op}$. Furthermore, for $\alpha\in C(Y_{0})$ and $\phi\in Y_{0}$,
$P(\mathcal{D}_{\alpha})(\phi)=\mathcal{D}_{\alpha}[\phi](\operatorname{id}_{G})=\alpha(\phi),$
and for $\mathcal{D}\in\mathsf{Op}$,
$\mathcal{D}_{P(\mathcal{D})}[\phi](h)=P(\mathcal{D})(\phi\cdot h)=\mathcal{D}[\phi\cdot h](\operatorname{id}_{G})=\mathcal{D}[\phi](h),$
so $\mathcal{D}_{P(\mathcal{D})}=\mathcal{D}$. ∎
Let $C^{\infty,\infty}(Y_{0}):=P(\mathsf{Op}^{\infty,\infty})$. Since elements of $C^{\infty,\infty}(Y_{0})$ correspond to operators from $Y_{0}$ to $C^{\infty,\infty}(G)$, we introduce some notation. For $\alpha\in C(Y_{0})$ and $\phi\in Y_{0}$, we let $\alpha_{\phi}:=\mathcal{D}_{\alpha}[\phi]\in C(G),$ i.e.,
$\alpha_{\phi}(g):=\alpha(\phi\cdot g)\text{\qquad for all }g\in G.$ (13)
Then by (12)
$\alpha_{\phi\cdot g}(h)=\alpha_{\phi}(gh)\text{\qquad for all }g,h\in G.$ (14)
By the following proposition, $C^{\infty,\infty}(Y_{0})$ is an algebra.
Proposition 4.3.
The following properties hold:
  1. (1)
    $C^{\infty,\infty}(Y_{0})$ is an algebra over $\mathbb{C}$, i.e., it is closed under addition, multiplication, and scalar multiplication.
  2. (2)
    For every $\alpha\in C^{\infty,\infty}(Y_{0})$ and every $X\in\mathfrak{g}$, there is an element $X\alpha\in C^{\infty,\infty}(Y_{0})$ such that for any $\phi\in Y_{0}$,
    $(X\alpha)_{\phi}=X[\alpha_{\phi}].$ (15)
    Furthermore, for any $X,Y\in\mathfrak{g}$ and $\alpha\in C^{\infty,\infty}(Y_{0})$,
    $[X,Y]\alpha=X[Y\alpha]-Y[X\alpha],$ (16)
    so $C^{\infty,\infty}(Y_{0})$ is a $\mathfrak{g}$–module.
  3. (3)
    For every $\alpha\in C(Y_{0})$, there is a sequence $(\alpha_{n})_{n}\in C^{\infty,\infty}(Y_{0})$ such that for any compact $K\subset Y_{0}$,
    $\lim_{n\to\infty}\|\alpha_{n}-\alpha\|_{L^{\infty}(K)}=0.$
Proof.
Because of the bijection between $C^{\infty,\infty}(Y_{0})$ and $\mathsf{Op}^{\infty,\infty}$, we can prove the proposition using either set.
For item (1), note that for $\alpha,\beta\in C(Y_{0})$, every $y\in Y_{0}$, and every $z\in\mathbb{C}$ we have $(\alpha+\beta)_{y}=\alpha_{y}+\beta_{y}$, $(\alpha\beta)_{y}=\alpha_{y}\beta_{y}$, and $(z\alpha)_{y}=z\alpha_{y}$, so $C^{\infty,\infty}(Y_{0})$ is closed under addition, multiplication, and scalar multiplication.
To prove item (2), for any left-invariant vector field $X$, $\mathcal{D}\in\mathsf{Op}^{\infty,\infty}$, and $\phi\in Y_{0}$, let
$X\mathcal{D}[\phi]:=X[\mathcal{D}[\phi]].$
It is straightforward to check that $X\mathcal{D}\in\mathsf{Op}^{\infty,\infty}$.
Furthermore, for $X,Y\in\mathfrak{g}$, $\mathcal{D}\in\mathsf{Op}^{\infty,\infty}$, and $\phi\in Y_{0}$,
$[X,Y]\mathcal{D}[\phi]=[X,Y][\mathcal{D}[\phi]]=XY[\mathcal{D}[\phi]]-YX[\mathcal{D}[\phi]]=X[Y\mathcal{D}][\phi]-Y[X\mathcal{D}][\phi],$
proving (16).
Finally, we prove item (3). For $h\in C(G)$ and $k\in C_{c}^{\infty}(G)$, let $h\ast k\in C^{\infty}(G)$,
$(h\ast k)(x):=\int_{G}h(y)k(y^{-1}x)\,\mathrm{d}y,$ (17)
Let $\mathcal{D}\in\mathsf{Op}$ and $k\in C_{c}^{\infty}(G)$. For $\phi\in Y_{0}$, let
$(\mathcal{D}\ast k)[\phi]=\mathcal{D}[\phi]\ast k\in C^{\infty,\infty}(G).$
It is readily verified that $\mathcal{D}\ast k$ is $G$–equivariant, so $\mathcal{D}\ast k\in\mathsf{Op}^{\infty,\infty}$. Thus, by Lemma 4.2, for any $\alpha\in C(Y_{0})$, we can define $\alpha\ast k\in C^{\infty,\infty}(Y_{0})$ as $\alpha\ast k=P(\mathcal{D}\ast k)$, so that $\mathcal{D}_{\alpha\ast k}=\mathcal{D}\ast k$.
Let $(k_{n})_{n}\in C_{c}^{\infty}(G)$ be a sequence of nonnegative functions such that
$\int_{G}k_{n}(x)\,\mathrm{d}x=1,$
and $\operatorname{supp}(k_{n})\subset B_{\frac{1}{n}}(\operatorname{id}_{G})$ for all $n$, so that for any $h\in C(G)$, $h\ast k_{n}\to h$ uniformly on compact sets as $n\to\infty$.
Let $\alpha_{n}:=\alpha\ast k_{n}$, so that for any $\phi\in Y_{0}$, we have
$\alpha_{n}(\phi)=(\alpha_{\phi}\ast k_{n})(\operatorname{id}_{G}).$
and thus $\lim_{n\to\infty}\alpha_{n}(\phi)=\alpha_{\phi}(\operatorname{id}_{G})=\alpha(\phi)$, i.e., $\alpha_{n}$ converges to $\alpha$ pointwise.
We claim that $\alpha_{n}$ converges to $\alpha$ uniformly on compact sets. Let $K\subset Y_{0}$ be a compact set. The function $(\phi,g)\mapsto\alpha_{\phi}(g)$ is continuous, so it is uniformly continuous on $K\times\overline{B}_{1}(\operatorname{id}_{G})$. That is, the functions $\alpha_{\phi},\phi\in K$, are uniformly continuous with a uniform modulus of continuity. Thus, for any $\varepsilon>0$, there is an $N>0$ such that for all $\phi\in K$ and $n>N$,
$|\alpha_{n}(\phi)-\alpha(\phi)|=|(\alpha_{\phi}\ast k_{n})(\operatorname{id}_{G})-\alpha_{\phi}(\operatorname{id}_{G})|<\varepsilon,$
i.e., $\|\alpha_{n}-\alpha\|_{L^{\infty}(K)}\to 0$ as $n\to\infty$. This proves item (3). ∎
4.2. Differential forms with twisted coefficients and the universal pullback map
Next, we define a complex of differential forms with coefficients in $C^{\infty,\infty}(Y_{0})$. Let $A=C^{\infty,\infty}(Y_{0})$, and for $k\geq 0$ let
$C^{k}(\mathfrak{g};A):=\mathrm{Hom}(\wedge^{k}\mathfrak{g},A).$
Any cochain $\beta\in C^{k}(\mathfrak{g};A)=\mathrm{Hom}(\wedge^{k}\mathfrak{g},A)$ corresponds to a function $\wedge^{k}\mathfrak{g}\times Y_{0}\to\mathbb{C}$. If we view $\beta$ instead as a function $Y_{0}\to\mathrm{Hom}(\wedge^{k}\mathfrak{g},\mathbb{C})$, we obtain a map $V_{\beta}\colon Y_{0}\to\wedge^{k}\mathfrak{g}^{*}$ such that $V_{\beta}(\phi)(\lambda)=\beta(\lambda)(\phi)$ for all $\phi\in Y_{0}$ and $\lambda\in\wedge^{k}\mathfrak{g}$. As before, we can define an operator $\mathcal{D}_{\beta}\colon Y_{0}\to\Omega^{k}(G)$ by letting $\mathcal{D}_{\beta}[\phi]$ be the $k$–form whose value at any $g\in G$ is
$\mathcal{D}_{\beta}[\phi]_{g}:=V_{\beta}(\phi\cdot g).$
Then for any left-invariant vector fields $v_{1},\dots,v_{k}\in\mathfrak{g}$,
$\mathcal{D}_{\beta}[\phi](v_{1},\dots,v_{k})=\mathcal{D}_{\beta(v_{1},\dots,v_{k})}[\phi].$ (18)
For conciseness, let $\beta_{\phi}:=\mathcal{D}_{\beta}[\phi]$.
Since $A$ is a $\mathfrak{g}$–module, we can define a differential and wedge product on $C^{*}(\mathfrak{g};A)$ by a standard construction: for every $f\in C^{k}(\mathfrak{g};A)$, and every $X_{1},\dots,X_{k+1}\in\mathfrak{g}$, let
$\mathrm{d}f(X_{1},\dots,X_{k+1}):=\sum_{i=1}^{k+1}(-1)^{i+1}X_{i}\left[f(X_{1},\dots,\hat{X}_{i},\dots,X_{k+1})\right]\\ +\sum_{i<j}(-1)^{i+j}f([X_{i},X_{j}],\dots,\hat{X}_{i},\dots,\hat{X}_{j},\dots,X_{k+1}),$ (19)
and for every $\alpha\in C^{m}(\mathfrak{g};A),\beta\in C^{n}(\mathfrak{g};A)$, let
$\alpha\wedge\beta:=\frac{(m+n)!}{m!n!}\mathcal{A}(\alpha\otimes\beta)\in C^{m+n}(\mathfrak{g};A),$ (20)
where $\alpha\otimes\beta$ is the tensor product and $\mathcal{A}$ is projection on alternating forms, i.e., for every $(0,k)$-tensor $t$, and every $k$-tuple $X_{1},\ldots,X_{k}\in\mathfrak{g}$
$\mathcal{A}t(X_{1},\ldots,X_{k}):=\frac{1}{k!}\sum_{\sigma\in S_{k}}\operatorname{sgn}(\sigma)t(X_{\sigma(1)},\ldots,X_{\sigma(k)}),$
where $\operatorname{sgn}(\sigma)$ is the signature of $\sigma$.
It is straightforward to check that for any $\alpha,\beta\in C^{*}(\mathfrak{g};A)$, and $\phi\in Y_{0}$,
$(\mathrm{d}\alpha)_{\phi}=\mathrm{d}[\alpha_{\phi}],$ (21)
and
$(\alpha\wedge\beta)_{\phi}=\alpha_{\phi}\wedge\beta_{\phi}.$ (22)
Consequently, one can verify that $C^{*}(\mathfrak{g};A)$ is a differential graded algebra either directly or by noting that
$\mathrm{d}[\mathrm{d}\alpha]_{\phi}=\mathrm{d}[[\mathrm{d}\alpha]_{\phi}]=\mathrm{d}[\mathrm{d}[\alpha_{\phi}]]=0,$
and
$\mathrm{d}[\alpha\wedge\beta]_{\phi}=(\mathrm{d}\alpha\wedge\beta+(-1)^{|\alpha|}\alpha\wedge\mathrm{d}\beta)_{\phi},$ (23)
for all $\phi$.
We use this structure to construct the map $T^{\sharp}\colon C^{*}(\mathfrak{h})\to C^{*}(\mathfrak{g};A)$ and prove Lemma 4.1.
Proof of Lemma 4.1.
For every $\omega\in C^{k}(\mathfrak{h})$ and $\lambda\in\wedge^{k}\mathfrak{g}$, let
$\mathcal{D}_{\omega,\lambda}[\phi]:=\phi^{*}\omega(\lambda)\in C^{\infty,\infty}(G).$
This operator is continuous and satisfies (10), so $\mathcal{D}_{\omega,\lambda}\in\mathsf{Op}^{\infty,\infty}$. By Lemma 4.2, there is a $T^{\sharp}\omega(\lambda)\in A$ such that
[…]
for all $\phi\in Y_{0}$, i.e., $T^{\sharp}$ satisfies (9). Furthermore, since $\phi^{*}\omega(\lambda)$ is linear in $\lambda$, $T^{\sharp}\omega(\lambda)$ is linear in $\lambda$ as well.
In order to show that $T^{\sharp}$ is a morphism of differential graded algebra, it remains to check that $T^{\sharp}$ commutes with the wedge product and differential. For every $\omega,\eta\in C^{k}(\mathfrak{h})$ and $\phi\in Y_{0}$, we have
[…]
Hence $T^{\sharp}(\omega\wedge\eta)=T^{\sharp}\omega\wedge T^{\sharp}\eta$.
Similarly, for every $\omega\in C^{k}(\mathfrak{h})$, and every $\phi\in Y_{0}$,
[…]
thus the proof is complete. ∎
5. Amenable averages and the vanishing lemma
In the previous sections, we showed that any ergodic map $\psi\in Y_{0}$ has a limit measure $\mu$ on $Y_{0}$ and we constructed a universal pullback map $T^{\sharp}\colon H^{*}(\mathfrak{h})\to H^{*}(\mathfrak{g};A)$, where $A=C^{\infty,\infty}(Y_{0})$ is a space of continuous functions on $Y_{0}$ corresponding to smooth $G$-equivariant operators $Y_{0}\to C^{\infty,\infty}(G)$. In this section, we show that elements of $C^{*}(\mathfrak{g};A)$ whose amenable average is zero are reduced coboundaries (i.e., limits of coboundaries) in a larger space $C^{*}(\mathfrak{g};L)$.
We first define the amenable average of a cochain. For any bounded form $\omega\in\Omega^{k}(G)$ and $p\in G$, denote the value of $\omega$ at $p$ by $\omega_{p}\in\wedge^{k}\mathfrak{g}^{*}$. We define the amenable average $\overline{\omega}\in\wedge^{k}\mathfrak{g}^{*}$ by
$\overline{\omega}=\lim_{n\to\infty}\fint_{B_{n}}\omega_{g}\,\mathrm{d}g,$ (24)
if the limit exists.
In the rest of this paper, we fix an ergodic map $\psi\in Y_{0}$ and let $\mu$ be its limit measure. When $\omega=\beta_{\psi}$, we can describe $\overline{\omega}$ in terms of $\mu$.
Lemma 5.1.
Let $\beta\in C^{k}(\mathfrak{g};A)$. Let $V_{\beta}\colon Y_{0}\to\wedge^{k}\mathfrak{g}^{*}$ as in Section 4.2. Then
$\overline{\beta_{\psi}}=\int V_{\beta}\,\mathrm{d}\mu.$
Proof.
As in Definition 3.1, for $n\geq 0$, let $\mu_{n}$ be the measure on $Y_{0}$ such that
$\int\alpha\,\mathrm{d}\mu_{n}=\fint_{B_{n}}\alpha(\psi\cdot g)\,\mathrm{d}g$
for any continuous $\alpha\in C(Y_{0})$, and $\mu_{n}$ converges weakly to $\mu$. In particular, since $V_{\beta}\colon Y_{0}\to\wedge^{k}\mathfrak{g}^{*}$ is continuous,
$\lim_{n\to\infty}\fint_{B_{n}}(\beta_{\psi})_{g}\,\mathrm{d}g=\lim_{n\to\infty}\fint_{B_{n}}V_{\beta}(\psi\cdot g)\,\mathrm{d}g=\lim_{n\to\infty}\int V_{\beta}\,\mathrm{d}\mu_{n}=\int V_{\beta}\,\mathrm{d}\mu.\qed$
In particular, $\overline{\beta_{\psi}}$ is determined by the values of $V_{\beta}$ on $\operatorname{supp}(\mu)$, which encourages us to study the space $L^{2}(\mu)$.
We equip $L^{2}(\mu)$ with the left action $g\cdot f(\phi)=f(\phi\cdot g)$ for all $\phi\in Y_{0}$ and $g\in G$. For any $v\in L^{2}(\mu)$, let $\eta_{v}\colon G\to L^{2}(\mu)$, $\eta_{v}(g)=g\cdot v$, and let
$(L^{2}(\mu))^{\infty}:=\{v\in L^{2}(\mu):\eta_{v}\text{ is smooth}\};$ (25)
let $L:=(L^{2}(\mu))^{\infty}$. We put a $\mathfrak{g}$–module structure on $L$ by letting
$Xv:=X\eta_{v}(\operatorname{id}_{G})=\frac{\mathrm{d}}{\mathrm{d}t}[\exp(tX)\cdot v]_{t=0}\in L,$ (26)
for all $X\in\mathfrak{g}$, and $v\in L$. Using this structure, we define the cochain complex $C^{*}(\mathfrak{g};L)$ in the usual way (see Section 4.2).
We equip $L$ with the topology such that $\lim_{n\to\infty}v_{n}=v$ if and only if
$\lim_{n\to\infty}\eta_{v_{n}}=\eta_{v}$
in the $C^{\infty}$ topology.
For $a\in A$ and $v\in L$, the function $a$ is bounded on $\operatorname{supp}(\mu)$, so the product $av$ lies in $L^{2}(\mu)$. Furthermore, one can check that this product satisfies the usual product rule
$X[av]=X[a]v+aX[v],$ (27)
for any $X\in\mathfrak{g}$, so in fact $av\in L$ and the map $(\alpha,v)\mapsto\alpha v$ is continuous.
Let
$L^{2}_{0}(\mu):=\left\{f\in L^{2}(\mu):\int_{Y_{0}}f\,\mathrm{d}\mu=0\right\},$ (28)
and let $L_{0}:=L\cap L^{2}_{0}(\mu)$. Since $\operatorname{supp}\mu$ is compact, every $\alpha\in A$ satisfies $\|\alpha\|_{L^{2}(\mu)}<\infty$. By Lemma 5.1, $\alpha\in L_{0}$ if and only if $\overline{\alpha_{\psi}}=0$. We equip $L_{0}$ with the topology it inherits as a subspace of $L$.
Our main goal in this section is to use results of van Est and Blanc to prove the following lemma.
Lemma 5.2.
For any $n\geq 0$,
$\overline{H}^{n}(\mathfrak{g};L_{0})=0.$
Thus, with $\psi$ and $\mu$ as above, if $\beta\in C^{n}(\mathfrak{g};A)$ is a cocycle and $\overline{\beta_{\psi}}=0$, then there is a sequence $\eta_{i}\in C^{n-1}(\mathfrak{g};L_{0})$, $i\in\mathbb{N}$, such that $\lim_{i}\mathrm{d}\eta_{i}=\beta$. That is, for any $\lambda\in\wedge^{n}\mathfrak{g}$, $\mathrm{d}\eta_{i}(\lambda)$ converges in $L$ to $\beta(\lambda)$.
We first recall some definitions. For any $G$–module $M$ and any $n\geq 0$, let
$F^{n}_{s}(G;M):=\{f:G^{n+1}\to M:\text{$f$ is smooth}\}.$ (29)
Let $\mathrm{d}=\mathrm{d}_{n}\colon F^{n}_{s}(G;M)\to F^{n+1}_{s}(G;M)$ be the coboundary operator
$\mathrm{d}f(g_{0},\dots,g_{n+1})=\sum_{i=0}^{n+1}(-1)^{i}f(g_{0},\dots,\hat{g}_{i},\dots,g_{n+1}),$
for $f\in F^{n}_{s}(G;M)$. Let $C_{s}^{n}(G;M)$ be the set of $G$–equivariant elements of $F^{n}_{s}(G;M)$, i.e., those that satisfy
$f(hg_{0},\ldots,hg_{n}):=h\cdot\left(f(g_{0},\ldots,g_{n})\right),\quad\forall h,g_{0},\ldots,g_{n}\in G.$
We equip $C^{n}_{s}(G;M)$ with the topology of $C^{\infty}$ convergence on compact subsets of $G^{n+1}$, and we define the smooth cohomology $H^{*}_{s}(G;M)$ of $G$ to be the cohomology of $C_{s}^{*}(G;M)$.
Blanc gave a condition for the reduced cohomology $\overline{H}_{s}^{*}(G;M)$ to vanish.
Theorem 5.3 ([Blanc]).
Let $G$ be a connected nilpotent Lie group and let $U$ be a unitary representation of $G$ with no nontrivial fixed points, i.e., if $v\in U$ and $g\cdot v=v$ for all $g$, then $v=0$. Let
$\overline{H}^{k}_{s}(G;U)=\ker\mathrm{d}_{k}/\operatorname{clos}(\operatorname{im}\mathrm{d}_{k-1}).$
Then $\overline{H}^{*}_{s}(G;U)=0$.
Blanc proved this theorem for $\overline{H}^{*}_{c}(G;U)$, the reduced continuous cohomology of $G$, but by a convolution argument (see [Guichardet, Equation (1.16) and Proposition 1.7, page 183]), there is a topological isomorphism between $\overline{H}^{*}_{c}(G;U)$ and $\overline{H}^{*}_{s}(G;U)$. The analogous statement when $G$ is a discrete group is Theorem 4.1.3 of [Shalom].
Suppose in addition that the map $g\mapsto g\cdot v$ is smooth for every $v\in M$ and $M$ is equipped with the $\mathfrak{g}$–module structure given by (26) (for instance, $M=L$ or $M=L_{0}$). In this case, van Est gave an isomorphism between $H_{s}^{*}(G;M)$ and $H^{*}(\mathfrak{g};M)$. (See [Guichardet, Section 7.3] for additional details on Theorem 5.4.)
Theorem 5.4 ([vanEstResultPrecise, Theorem 14.1], [Guichardet, Corollaire 7.2]).
Let $M$ be as above. For all $n$, let $R\colon C_{s}^{*}(G;M)\to C^{*}(\mathfrak{g};M)$ be the chain map such that for $f\in C_{s}^{n}(G;M)$,
$Rf(X_{1},\dots,X_{n}):=\sum_{\sigma\in S_{n}}\operatorname{sgn}(\sigma)X_{\sigma_{1},2}\dots X_{\sigma_{n},n+1}f(\operatorname{id}_{G},\dots,\operatorname{id}_{G}),$
where $X_{i,j}f$ denotes the derivative along $X_{i}$ in the $j$th argument of $f$.
Then $R$ induces an isomorphism from $H_{s}^{*}(G;M)$ to $H^{*}(\mathfrak{g};M)$.
We use these results to prove Lemma 5.2.
Proof of Lemma 5.2.
In the following, we equip $L^{2}(\mu)$ and $L^{2}_{0}(\mu)$ with the topology induced by the $L^{2}$ norm, and we equip $L_{0}$ with the $C^{\infty}$ topology induced by the inclusion $L_{0}\subset C^{\infty}(G,L^{2}_{0}(\mu))$, $v\mapsto\eta_{v}$. We view $C^{d}_{s}(G;V)$ as a space of functions from $G^{d+1}$ to $V$ and equip it with the $C^{\infty}$ topology.
For $k$ a compactly supported smooth function on $G$, $V$ a Hilbert space, and $f\colon G\to V$, let
$(f\ast k)(x):=\int_{G}f(xg)k(g^{-1})\,\mathrm{d}g,$
for any $x\in G$. For $v\in L^{2}_{0}(\mu)$, let $v\ast k\in L^{2}_{0}(\mu)$,
$v\ast k:=\int_{G}(g\cdot v)k(g^{-1})\,\mathrm{d}g,$
so that for $x\in G$,
$\eta_{v\ast k}(x)=x\cdot\int_{G}(g\cdot v)k(g^{-1})\,\mathrm{d}g=\int_{G}(xg\cdot v)k(g^{-1})\,\mathrm{d}g=(\eta_{v}\ast k)(x).$
Then for $h\in G$ and $X\in\mathfrak{g}$,
$X\eta_{v\ast k}(h)$ $=\frac{\mathrm{d}}{\mathrm{d}t}\bigg{|}_{t=0}\int_{G}(h\exp(tX)g\cdot v)k(g^{-1})\,\mathrm{d}g$
$=\frac{\mathrm{d}}{\mathrm{d}t}\bigg{|}_{t=0}\int_{G}(hg\cdot v)k(g^{-1}\exp(tX))\,\mathrm{d}g=\eta_{v\ast Xk}(h).$
Since $k$ is smooth, $v\ast k\in L_{0}$. Furthermore, if $v_{i},w\in L^{2}_{0}(\mu)$, and $\|v_{i}-w\|_{L^{2}(\mu)}\to 0$, then $\eta_{v_{i}\ast k}$ converges to $\eta_{w\ast k}$ in the $C^{\infty}$ topology. That is, for any $k$, the map $v\mapsto v\ast k$ is a continuous linear map from $L^{2}_{0}(\mu)$ to $L_{0}$.
For $k$ as above and $\sigma\in C^{d}_{s}(G;L^{2}_{0}(\mu))$, we let $\sigma\ast k\in C^{d}_{s}(G;L_{0})$,
$(\sigma\ast k)(g_{0},\dots,g_{d}):=\sigma(g_{0},\dots,g_{d})\ast k,$
for all $g_{0},\dots,g_{d}\in G$. This satisfies
$\mathrm{d}[v\ast k]=\mathrm{d}[v]\ast k,$
and $\sigma\mapsto\sigma\ast k$ is a continuous linear map from $C_{s}^{d}(G;L^{2}_{0}(\mu))$ to $C_{s}^{d}(G;L_{0})$.
Now we prove the lemma. Let $R\colon C_{s}^{*}(G;L_{0})\to C^{*}(\mathfrak{g};L_{0})$ be the chain map from Theorem 5.4. Note that $R$ is continuous. Let $S\colon C^{*}(\mathfrak{g};L_{0})\to C_{s}^{*}(G;L_{0})$ be a homotopy inverse of $R$. Let $\beta\in C^{d}(\mathfrak{g};L_{0})$ such that $\mathrm{d}\beta=0$ and let $\gamma\in C^{d-1}(\mathfrak{g};L_{0})$ be such that $\mathrm{d}\gamma=\beta-RS\beta$.
Since $\mu$ is ergodic, $L^{2}_{0}(\mu)$ satisfies the assumption of Theorem 5.3. Therefore, there are $h_{1},h_{2},\dots\in C^{d-1}_{s}(G;L^{2}_{0}(\mu))$ such that $\mathrm{d}h_{n}$ converges to $S\beta$ in $C^{d}_{s}(G;L^{2}_{0}(\mu)).$
Let $k_{1},k_{2},\dots\colon G\to\mathbb{R}$ be smooth nonnegative functions such that $\operatorname{supp}k_{i}\subset B_{\frac{1}{i}}(\operatorname{id}_{G})$ and $\int_{G}k_{i}(x)\,\mathrm{d}x=1$ for all $i$, so that $\lim_{i}c\ast k_{i}=c$ for any $c\in C^{d}_{s}(G;L_{0})$.
Then $\mathrm{d}[h_{n}\ast k_{i}]=\mathrm{d}[h_{n}]\ast k_{i}$, and since the convolution by $k_{i}$ is continuous,
$\lim_{n}\mathrm{d}[h_{n}]\ast k_{i}=S\beta\ast k_{i}$
as elements of $C^{d}_{s}(G;L_{0})$. Since $\lim_{i}S\beta\ast k_{i}=S\beta$, we can choose a sequence $n_{i}\geq 0$ such that
$\lim_{i}\mathrm{d}[h_{n_{i}}\ast k_{i}]=S\beta.$
Let $\hat{h}_{i}:=h_{n_{i}}\ast k_{i}$. By the continuity of $R$,
$\lim_{i\to\infty}\,\mathrm{d}[\gamma+R\hat{h}_{i}]=(\beta-RS\beta)+RS\beta=\beta,$
so $\overline{H}^{d}(\mathfrak{g};L_{0})=0$.
Finally, if $\beta\in C^{n}(\mathfrak{g};A)$ is a cocycle and $\overline{\beta_{\psi}}=0$, then by Lemma 5.1, $\int V_{\beta}\,\mathrm{d}\mu=\overline{\beta_{\psi}}=0$. In particular, for any $\lambda\in\wedge^{n}\mathfrak{g}$,
$\int\beta(\lambda)(\phi)\,\mathrm{d}\mu(\phi)=\int V_{\beta}(\phi)(\lambda)\,\mathrm{d}\mu=\overline{\beta_{\psi}}(\lambda)=0,$
so $\beta(\lambda)\in L_{0}$ and thus $\beta\in C^{n}(\mathfrak{g};L_{0})$. Therefore, the reduced cohomology class of $\beta$ is zero, as desired. ∎
Remark 5.5.
Let $A_{0}:=\{f\in A\mid\int_{Y_{0}}f\,\mathrm{d}\mu=0\}$. Then Lemma 5.2 shows that any cocycle $\beta\in C^{d}(\mathfrak{g};A_{0})$ vanishes in $\overline{H}^{d}(\mathfrak{g};L_{0})$; it is unclear whether $\beta$ also vanishes in $\overline{H}^{d}(\mathfrak{g};A_{0})$, and the relationship between $A_{0}$ and $L_{0}$ is subtle.
On one hand, we can relate the $L^{2}(\mu)$ norm to a norm on certain functions on $G$. For $\gamma\colon G\to\mathbb{C}$, we define the amenable norm of $\gamma$ by
$\|\gamma\|_{M}:=\sqrt{\lim_{n\to\infty}\fint_{B_{n}}|\gamma(g)|^{2}\,\mathrm{d}g},$ (30)
if the limit exists. If $\beta\in A$ and $\psi\in Y_{0}$ is an ergodic map with limit measure $\mu$, then
$\|\beta_{\psi}\|_{M}=\sqrt{\lim_{n\to\infty}\fint_{B_{n}}|\beta(\psi\cdot g)|^{2}\,\mathrm{d}g}=\sqrt{\int|\beta(y)|^{2}\,\mathrm{d}\mu(y)}=\|\beta\|_{L^{2}(\mu)}.$
Thus, if $\beta\in A$ and $\|\beta\|_{L^{2}(\mu)}$ is small, then $\beta_{\psi}$ has small amenable norm.
By Lusin’s Theorem and Proposition 4.3.(3), any $f\in L$ is the limit in $L^{2}(\mu)$ of a sequence of functions $\alpha_{i}\in A$; when $i$ and $j$ are large, $\|\alpha_{i}-\alpha_{j}\|_{L^{2}(\mu)}$ is small, so $\|(\alpha_{i})_{\psi}-(\alpha_{j})_{\psi}\|_{M}$ is small.
On the other hand, while we can approximate any $f\in L$ in the $L^{2}(\mu)$ norm by functions $\alpha_{i}\in A$, it is unclear whether we can approximate $f$ in the topology on $L$. In particular, for $X\in\mathfrak{g}$, the derivative operator $h\mapsto X[h]$ from $L$ to $L$ is unbounded in the $L^{2}(\mu)$ norm, so it is not clear whether we can choose the $\alpha_{i}$’s such that the derivatives $X[\alpha_{i}]$ also converge in $L^{2}(\mu)$. It is possible that $L$ is substantially larger than $A$, but we have not been able to construct explicit elements of $L$ that do not lie in the closure of $A$.
6. Amenable averages and cohomology
In this section, we will use the maps constructed in the previous sections to prove the first part of Theorem 1.1. That is, we will show that if $\psi\in Y_{0}$ is an ergodic map, then the map $\overline{\psi}\colon H^{*}(\mathfrak{h})\to H^{*}(\mathfrak{g})$ defined by (2) induces a homomorphism of cohomology algebras.
We first write $\overline{\psi}$ in terms of a projection from $L$ to $\mathbb{C}$. As in the previous sections, let $\psi\in Y_{0}$ be an ergodic map, let $\mu$ be its limit measure, and let $L=(L^{2}(\mu))^{\infty}$ defined in (25). Let $M\colon L\to\mathbb{C}$,
$Mf:=\int_{Y_{0}}f\,\mathrm{d}\mu.$
Since $\mu$ is $G$–invariant, $M$ is a continuous $G$–module homomorphism with respect to the trivial $G$–action on $\mathbb{C}$. Therefore, $M$ induces maps from $C^{*}(\mathfrak{g};L)$ to $C^{*}(\mathfrak{g})$, and from $H^{*}(\mathfrak{g};L)$ to $H^{*}(\mathfrak{g})$. We denote these maps by $M$ as well.
When $f\in A$, Lemma 5.1 lets us write $Mf$ in terms of the amenable average:
$Mf=\int_{Y_{0}}f\,\mathrm{d}\mu=\lim_{n\to\infty}\fint_{B_{n}}f_{\psi}\,\mathrm{d}g=\overline{f_{\psi}}.$
Likewise, when $\alpha\in C^{d}(\mathfrak{g};A)$ and $\lambda\in\wedge^{d}\mathfrak{g}$,
[…]
Lemma 6.1.
Let $T^{\sharp}$ be the map defined in Lemma 4.1. For every left-invariant form $\omega\in C^{d}(\mathfrak{h})$, the amenable average
$\overline{\psi^{*}\omega}:=\lim_{n\to\infty}\fint_{B_{n}}(\psi^{*}\omega)_{g}\,\mathrm{d}g$
exists and satisfies $\overline{\psi^{*}\omega}=(M\circ T^{\sharp})(\omega)$.
Proof.
Let $\omega\in C^{d}(\mathfrak{h})$. Then $M(T^{\sharp}\omega)$ exists and for any $\lambda\in\wedge^{d}\mathfrak{g}$,
[…]
which proves the lemma. ∎
Let $\overline{\psi^{*}}:=M\circ T^{\sharp}\colon C^{*}(\mathfrak{h})\to C^{*}(\mathfrak{g})$. By Lemma 6.1, this satisfies (2). Since $T^{\sharp}$ is a homomorphism of differential graded algebras, $T^{\sharp}$ induces a homomorphism of cohomology algebras. We claim that $M$ also induces a homomorphism of cohomology algebras, which would imply that $\overline{\psi^{*}}$ does as well. For $u\in C^{*}(\mathfrak{g};A)$, let $[u]\in H^{*}(\mathfrak{g};A)$ denote the corresponding cohomology class.
Lemma 6.2.
The map $M$ induces a homomorphism of cohomology algebras from $H^{*}(\mathfrak{g};A)$ to $H^{*}(\mathfrak{g})$. That is, for every $[u_{1}],[u_{2}]\in H^{*}(\mathfrak{g};A)$,
$[Mu_{1}\wedge Mu_{2}]=[M(u_{1}\wedge u_{2})],\qquad\text{in $H^{*}(\mathfrak{g})$}.$ (32)
Proof.
Let $L:=(L^{2}(\mu))^{\infty}$, $L_{0}:=L^{2}_{0}(\mu)\cap L$, and $A_{0}:=\{f\in A:\int_{Y_{0}}f\,\mathrm{d}\mu=0\}$ as in the previous sections. We can decompose $A=A_{0}\oplus\mathbb{C}$ and $L=L_{0}\oplus\mathbb{C}$, where $\mathbb{C}$ is the space of constant functions on $Y_{0}$; then $M\colon L\to\mathbb{C}$ is the projection to $\mathbb{C}$.
For a module $Z$, let $\overline{B}^{k}(\mathfrak{g};Z):=\operatorname{clos}(\mathrm{d}^{k-1})\subset C^{k}(\mathfrak{g};Z))$ be the space of reduced coboundaries with coefficients in $Z$. Let $\alpha\in C^{m}(\mathfrak{g};A)$, $\beta\in C^{n}(\mathfrak{g};A)$ be cocycles. We claim that $M(\alpha)\wedge M(\beta)$ and $\alpha\wedge\beta$ lie in the same reduced cohomology class in $\overline{H}^{*}(\mathfrak{g};L)$, i.e.,
$M(\alpha)\wedge M(\beta)\equiv\alpha\wedge\beta\pmod{\overline{B}^{*}(\mathfrak{g};L)}.$ (33)
Let $\beta_{0}:=\beta-M\beta$. Then $\beta_{0}$ is a cocycle in $C^{n}(\mathfrak{g};A_{0})$. By Lemma 5.1, and Lemma 5.2, there is a sequence $(\eta_{i})_{i}\in C^{n-1}(\mathfrak{g};L_{0})$ such that $\lim_{i}\eta_{i}=\mathrm{d}\beta_{0}$.
Recall that for $a\in A$ and $b\in L$, we can define a continuous product $ab\in L$ that satisfies the usual product rule (27). Thus, for $\alpha\in C^{m}(\mathfrak{g};A)$ and $\eta_{i}\in C^{n-1}(\mathfrak{g};L_{0})$, we can define the wedge product $\alpha\wedge\eta_{i}$ as in (20). This is continuous, so
$\lim_{i\to\infty}\mathrm{d}[\alpha\wedge\eta_{i}]=\lim_{i\to\infty}\mathrm{d}[\alpha]\wedge\eta_{i}+(-1)^{m}\alpha\wedge\mathrm{d}[\eta_{i}]=(-1)^{m}\alpha\wedge\beta_{0}.$
That is, $\alpha\wedge(\beta-M\beta)\in\overline{B}^{*}(\mathfrak{g};L)$, so $\alpha\wedge\beta\equiv\alpha\wedge M(\beta)\pmod{\overline{B}^{*}(\mathfrak{g};L)}.$
Applying this argument to $M(\beta)$ and $\alpha$, we find $\alpha\wedge M(\beta)\equiv M(\alpha)\wedge M(\beta)$, so
$\alpha\wedge\beta\equiv\alpha\wedge M(\beta)\equiv M(\alpha)\wedge M(\beta)\pmod{\overline{B}^{*}(\mathfrak{g};L)}.$ (34)
We have $M(\overline{B}^{*}(\mathfrak{g};L))\subset\overline{B}^{*}(\mathfrak{g})=B^{*}(\mathfrak{g})$, where the last equality follows from the finite-dimensionality of $C^{*}(\mathfrak{g})$. Therefore, we can apply $M$ to both sides of (34) to conclude
$M(\alpha\wedge\beta)\equiv M(M(\alpha)\wedge M(\beta))=M(\alpha)\wedge M(\beta)\pmod{B^{*}(\mathfrak{g})},$
as desired. ∎
This proves the first part of Theorem 1.1.
7. Quasi-isometries and cohomology
Finally, we prove that if $\psi\in Y_{0}$ is an ergodic quasi-isometry, then $\overline{\psi^{*}}$ is an isomorphism of cohomology algebras. We first show that $\overline{\psi^{*}}$ preserves the top-dimensional cohomology.
We will need the following lemma. Recall that, for $C>0$, and $L\geq 1$, $\psi:G\to H$ is an $(L,C)$–quasi-isometry if
$L^{-1}d_{G}(g_{1},g_{2})-C\leq d_{H}(\psi(g_{1}),\psi(g_{2}))\leq Ld_{G}(g_{1},g_{2})+C,\quad\forall g_{1},g_{2}\in G,$ (35)
and if for all $h\in H$, there is a $g\in G$ such that $d_{H}(\psi(g),h)\leq C$. We say that $\psi$ is an $(L,C)$-quasi-isometric embedding if we only have (35).
Lemma 7.1.
Let $G$ and $H$ be two simply connected nilpotent Lie groups. Fix orientations on $G$ and $H$ and let $\psi\colon G\to H$ be a Lipschitz $(L,C)$–quasi-isometry. Let $\theta$ be the degree of $\psi$, let $\omega$ be the volume form on $H$, and let
$\tau(R):=\int_{B_{R}^{G}}\psi^{*}\omega\,\mathrm{d}g.$
Then $\theta\in\{-1,1\}$, and there is a $c$ depending on $G$, $H$, $L$, and $C$ such that $\theta\tau(R)\geq c|B_{R}^{G}|$ for all sufficiently large $R$.
Given this lemma, we can show that $\overline{\psi^{*}}([\omega])\neq[0]$.
Corollary 7.2.
Let $\psi\in Y_{0}$ be an ergodic quasi-isometry, and $\omega$ be the volume form on $H$. Let $n$ be the topological dimension of $H$. Then $[\omega]$ is a generator of $H^{n}(\mathfrak{h})$, and
$\overline{\psi^{*}}([\omega])\neq[0].$ (36)
Proof.
Let $\lambda\in\wedge^{n}\mathfrak{g}$ be the positively oriented unit $n$–vector. Assume $\theta=1$ in Lemma 7.1, the other case is analogous. Then
$|\overline{\psi^{*}}(\omega)(\lambda)|=\lim_{R\to\infty}|B^{G}_{R}|^{-1}\tau(R)\geq c,$
so $\overline{\psi^{*}}(\omega)\neq 0$, and (36) follows because $G$ is $n$-dimensional too. ∎
Before proving Lemma 7.1, we recall some facts about the degree of a map and the area formula, which can be found in, for instance, [MappingDegreeTheory, Section IV], and [Federer1996GeometricTheory], respectively. We fix two simply connected nilpotent Lie groups $G$ and $H$ of the same topological dimension, and equip them with Riemannian metrics and orientations. We recall that a map $f\colon X\to Y$ between topological spaces is proper if $f^{-1}(y)$ is compact for every $y\in Y$.
Let $U\subset G$ be an open set with closure $\overline{U}$, let $f\colon\overline{U}\subset G\to H$ be a continuous function, and $h\in H$. We say that $(f,U,h)$ is admissible if $f$ is proper, and $h\not\in f(\partial U)$. If in addition $f$ is smooth, and $h$ is a regular value for $f$, we define
$\deg(f,U,h):=\sum_{x\in f^{-1}(h)\cap U}\operatorname{sgn}(Df_{x}),$ (37)
where $\operatorname{sgn}(Df_{x})=1$ if $Df_{x}$ is orientation-preserving and $-1$ if $Df_{x}$ is orientation-reversing. Since $f$ is proper and $h$ is a regular value, $f^{-1}(h)$ is a finite set and this is a finite sum.
When $(f,U,h)$ is admissible, but $f$ is not smooth or $h$ is not a regular value, we can still define $\deg(f,U,h)$ by approximating by smooth functions. We let
$\deg(f,U,h):=\deg(\tilde{f},U,h^{\prime}),$
where $(\tilde{f},U,h^{\prime})$ is admissible, $\tilde{f}\colon\overline{U}\to H$ is smooth, $h^{\prime}$ is a regular value for $\tilde{f}$, and
$d_{H}(h,h^{\prime})+\sup_{g\in\overline{U}}d_{H}(f(g),\tilde{f}(g))\ll\mathrm{dist}_{H}(h,f(\partial U)).$
(In particular, when $U=G$, we require that $\sup_{g\in G}d_{H}(f(g),\tilde{f}(g))<\infty$.)
It is well-known that this definition is independent of the choice of $\tilde{f}$ and $h^{\prime}$. In fact, one can show the following, see [MappingDegreeTheory, Proposition 2.3, page 146 & Proposition 2.4, page 147 & Corollary 2.5(3), page 149].
Proposition 7.3.
Let $U\subset G$ be an open set and let $f\colon\overline{U}\to H$ be a proper map.
  1. (1)
    (Basepoint invariance) If $h$ and $h^{\prime}$ lie in the same path component of $H\setminus f(\partial(U))$, then $\deg(f,U,h)=\deg(f,U,h^{\prime})$. Further, if $f\colon G\to H$ is proper, then $\deg(f,G,h)$ is independent of $h$, and we define $\deg(f):=\deg(f,G,h)$.
  2. (2)
    (Homotopy invariance) If $f_{t}\colon\overline{U}\to H,t\in[0,1]$ is a homotopy such that $f_{t}$ is proper and $h\not\in f_{t}(\partial U)$ for all $t$, then $\deg(f_{0},U,h)=\deg(f_{1},U,h)$.
  3. (3)
    (Excision) If $D\subset U$ is an open set and $f^{-1}(h)\subset D$, then $\deg(f,U,h)=\deg(f,D,h)$.
We also recall the area formula, see, e.g., [Federer1996GeometricTheory, Theorem 3.2.3], and [Federer1996GeometricTheory, Corollary 4.1.26]. Let $U\subset G$ be an open set, and let $f\colon\overline{U}\to H$ be a proper Lipschitz map. Then, denoting with $\omega$ the volume form on $H$, we have
$\int_{U}f^{*}\omega=\int_{U}\det(Df(g))\,\mathrm{d}g=\int_{f(U)}\deg(f,U,h)\,\mathrm{d}h,$ (38)
and
$\int_{U}|\det(Df(g))|\,\mathrm{d}g=\int_{H}\#f^{-1}(h)\,\mathrm{d}h\geq\int_{f(U)}|\deg(f,U,h)|\,\mathrm{d}h.$ (39)
Finally, recall that if $G$ is a simply connected nilpotent Lie group, then the rescalings $(G,r^{-1}d_{G})$ Gromov–Hausdorff converge, as $r\to\infty$, to a scaling limit $G^{\infty}$, which is the associated graded nilpotent Lie group equipped with a subRiemannian metric. Furthermore, there is a family of homeomorphisms $s^{G}_{r}\colon G^{\infty}\to G,r>0$ with $s^{G}_{r}(0)=0$ such that for all $v,v^{\prime}\in G^{\infty}$,
$\lim_{r\to\infty}r^{-1}d_{G}(s_{r}(v),s_{r}(v^{\prime}))=d_{G^{\infty}}(v,v^{\prime}),$
locally uniformly in $v$ and $v^{\prime}$, and such that for all $R>0$,
$\lim_{r\to\infty}r^{-1}d_{\text{Haus}}(s^{G}_{r}(B_{R}^{G^{\infty}}),B_{rR}^{G})=0.$
In fact, we can write $s^{G}_{r}=\exp_{G}\circ M_{r}\circ\exp_{G^{\infty}}^{-1}$, where $M_{r}$ is a family of linear maps from $\mathfrak{g}^{\infty}$ to $\mathfrak{g}$, see the scheme of the proof in [LeDonnePrimer, Theorem 4.7], and the original proof in [PansuCroissance, Section (38)–(49)], in particular [PansuCroissance, Proposition (47)].
If $\psi\colon G\to H$ is a quasi-isometry such that $\psi(0)=0$ and $s^{H}_{r}\colon H^{\infty}\to H,r>0$ is the corresponding family for $H$, then we can define a family of maps $\psi_{r}\colon G^{\infty}\to H^{\infty}$,
$\psi_{r}:=(s_{r}^{H})^{-1}\circ\psi\circ s_{r}^{G}.$ (40)
Then there is a subsequence $(r_{i})_{i}$, with $r_{i}\to\infty$, such that $\psi_{r_{i}}$ converges locally uniformly to a bilipschitz homeomorphism $\psi_{\infty}\colon G^{\infty}\to H^{\infty}$ as $i\to\infty$.
For $U\subset G$ or $U\subset H$, let $|U|$ be the Riemannian volume of $U$. For $\Lambda=G,H$ and $R>0$, let $B_{R}^{\Lambda}\subset\mathbb{R}^{n}$ be the unit ball centered at 00 with respect to $d_{\Lambda}$.
Proof of Lemma 7.1.
Without loss of generality, suppose that $\psi(0)=0$. For $r>0$, let $\psi_{r}\colon G^{\infty}\to H^{\infty}$ as in (40), and let $(r_{i})_{i}$ be a sequence such that $\psi_{r_{i}}$ converges locally uniformly to $\psi_{\infty}\colon G^{\infty}\to H^{\infty}$, as $r_{i}\to\infty$.
Suppose that $\psi_{\infty}$ is positively oriented, i.e., $\deg(\psi_{\infty})=1$; the negatively oriented case is similar. Let $B=B^{G^{\infty}}_{1}$ be the unit ball in $G^{\infty}$. Since $\psi$ is a quasi-isometry, it is proper, and $\psi^{-1}(0)$ is compact. Thus, when $i$ is sufficiently large, $s^{G}_{r_{i}}(B)\Supset\psi^{-1}(0)$, and by excision, $\deg(\psi,G,0)=\deg(\psi,s^{G}_{r_{i}}(B),0)$. Therefore,
$\deg(\psi)=\lim_{i\to\infty}\deg(\psi,s^{G}_{r_{i}}(B),0)=\lim_{i\to\infty}\deg(\psi_{r_{i}},B,(s_{r_{i}}^{H})^{-1}(0))=\lim_{i\to\infty}\deg(\psi_{r_{i}},B,0).$
Since $\psi_{r_{i}}$ converges uniformly to $\psi_{\infty}$ on $B$, the homotopy invariance of the degree implies that
$\lim_{i\to\infty}\deg(\psi_{r_{i}},B,0)=\deg(\psi_{\infty},B,0)=1,$
and thus $\deg(\psi)=\deg(\psi,G,0)=1$.
Let $R>4CL$ and let $R^{\prime}=R-2CL$. If $x\in B_{R^{\prime}}^{G}$ and $y\not\in B_{R}^{G}$, then $d_{G}(x,y)\geq 2CL$, so, since $\psi$ is an $(L,C)$-quasi-isometry, $d_{H}(\psi(x),\psi(y))>0$, i.e., $\psi(x)\neq\psi(y)$. Let $Z=\psi(B_{R}^{G})$ and $Z^{\prime}=\psi(B_{R^{\prime}}^{G})$. Then $\psi^{-1}(Z^{\prime})\subset B_{R}^{G}$, so by excision, for all $z\in Z^{\prime}$,
$\deg(\psi,B_{R}^{G},z)=\deg(\psi,G,z)=1.$ (41)
Now we calculate $\tau(R)=\int_{B_{R}^{G}}\psi^{*}\omega\,\mathrm{d}g$. By the area formula (38), and (41),
[…]
Then
$\big{|}\tau(R)-|Z^{\prime}|\big{|}\leq\int_{Z\setminus Z^{\prime}}|\deg(\psi,B_{R}^{G},z)|\,\mathrm{d}z\leq\int_{Z\setminus Z^{\prime}}\#\left(B_{R}^{G}\cap\psi^{-1}(z)\right)\,\mathrm{d}z.$
Let $M:=B_{R}^{G}\setminus B_{R^{\prime}}^{G}$. Then $B_{R}^{G}\cap\psi^{-1}(z)\subset M$ for every $z\in Z\setminus Z^{\prime}$, so, using (39),
$\big{|}\tau(R)-|Z^{\prime}|\big{|}\leq\int_{H}\#\left((\psi|_{M})^{-1}(z)\right)\,\mathrm{d}z=\int_{M}|\det(Df(g))|\,\mathrm{d}x\leq|M|\operatorname{Lip}(\psi)^{n}.$
If $d_{G}(0,g)>R^{\prime}$, then $d_{H}(0,\psi(g))>L^{-1}R^{\prime}-C\geq\frac{R}{4L}$, so $\psi^{-1}(B_{\frac{R}{4L}}^{H})\subset B_{R^{\prime}}^{G}$. Since $\psi$ is degree–$1$, it is surjective, so this implies $B_{\frac{R}{4L}}^{H}\subset\psi(B_{R^{\prime}}^{G})$. Thus,
$|Z^{\prime}|=\left|\psi(B_{R^{\prime}}^{G})\right|\geq\left|B_{\frac{R}{4L}}^{H}\right|.$
Since $G$ and $H$ are quasi-isometric, they have the same growth exponent $Q$, and there is a $c>0$ depending only on $G,H,L$, and $C$ such that when $R$ is large, $|Z^{\prime}|\geq cR^{Q}$. Likewise, there is a $c^{\prime}>0$ such that when $R$ is large,
$|M|\operatorname{Lip}(\psi)^{n}=\operatorname{Lip}(\psi)^{n}(\left|B_{R}^{G}\right|-\left|B_{R^{\prime}}^{G}\right|)\leq c^{\prime}R^{Q-1}.$
Thus, when $R$ is sufficiently large,
$\tau(R)\geq|Z^{\prime}|-|M|\operatorname{Lip}(\psi)^{n}\geq cR^{Q}-c^{\prime}R^{Q-1}\geq\frac{c}{2}R^{Q},$
as desired. ∎
Finally, we complete the proofs of Theorem 1.3 and Theorem 1.1.
Proof of Theorem 1.1.
As proved in Section 6, a direct application of Lemma 6.1, Lemma 4.1, and Lemma 6.2, gives that the map $\overline{\psi^{*}}$ in (2) exists, and induces a homomorphism $\overline{\psi^{*}}:H^{*}(\mathfrak{h})\to H^{*}(\mathfrak{g})$ of cohomology algebras over $\mathbb{C}$.
If $\psi$ is a quasi-isometry, then by Corollary 7.2 we have $\overline{\psi^{*}}([\omega])\neq[0]$, where $[\omega]$ is a generator of the top-dimensional cohomology group. Hence, by Poincaré duality, $\overline{\psi^{*}}$ is injective. Switching the roles of $\mathfrak{g}$ and $\mathfrak{h}$, there is also an injective algebra homomorphism $H^{*}(\mathfrak{g})\to H^{*}(\mathfrak{h})$. Hence, $H^{*}(\mathfrak{g})$ and $H^{*}(\mathfrak{h})$ have the same dimension, so $\overline{\psi^{*}}$ is an isomorphism of cohomology algebras over $\mathbb{C}$, as desired. ∎
Proof of Theorem 1.3.
Suppose that $G$ and $H$ are quasi-isometric simply connected nilpotent Lie groups. By [BoissonatDyerGhosh], $G$ admits a triangulation with bounded geometry, so if $f\colon G\to H$ is a quasi-isometry, there is a continuous quasi-isometry $f^{\prime}\colon G\to H$ which agrees with $f$ on the vertices of the triangulation. (One can also construct $f$ using the nerve of a collection of points in $G$, see for instance Chapter 3.8 of [DrutuKapovich].)
By a convolution argument (see for instance [KMX1, Lemma 3.33]), there exists a quasi-isometry $\phi$ with bounded derivatives, i.e., $\phi\in Y_{0}$. Then, by Proposition 1.2 there exists an ergodic quasi-isometry $\psi\in Y_{0}$. By Theorem 1.1, the pullback $\overline{\psi^{*}}$ induces an isomorphism of cohomology algebras $H^{*}(\mathfrak{h})\cong H^{*}(\mathfrak{g})$ as desired. ∎
[…]
[title=References]